OpenStudy (anonymous):

Write the solution in lowest terms. 5/8 - 3/4 =

OpenStudy (anonymous):

Multiply by the common denominator which is 8 and 8 goes into itself 1 and into four twice \[\frac{5*1}{8*1}-\frac{3*2}{4*2}\]

OpenStudy (theviper):

\[\LARGE{\color{red}{\frac{5}{8}-\frac{3}{4}=\frac{-1}{8}}}\]

OpenStudy (anonymous):

\[=\frac{5*1}{8*1}-\frac{3*2}{4*2}\] \[=\frac{5}{8}-\frac{6}{8}\] \[=-\frac{1}{8}\]